The relation (3t=3x+6) describes the displacement of a particle in one direction. Where x is in metres and t in seconds. Find the displacement when the velocity is zero


1
Expert's answer
2021-02-22T10:23:44-0500

The relation "3t=\sqrt{3}x+6"  describes the displacement of particle in one direction where x is in meter and t is in second.


We have to find displacement when velocity is zero. i.e. "\dfrac{dx}{dt}"

Now, "3t=\sqrt{3}t+6"


"3t-6=\sqrt{3}x"


"(3t-6)^2=3x"


"x = \dfrac{9t^2-36t+36}{3}"


Differentiate w.r.t 't'


"\dfrac{dx}{dt} = \dfrac{18t-36}{3} = 0"


"t=2s"


Hence, the displacement at "t = 2" is


"x = 0" m